Is the following statement true or false. Give an explanation and an example of your answer.


if f' is continuous on [1, 3], then the integral which goes from 1 to 3 "f' (v) d(v) = f(3) - f(1)" ?


1
Expert's answer
2022-04-18T00:18:39-0400

This statement is true, according to the fundamental theorem of calculus.

Let "f(x) = 2x^2". Then "f'(x) = 4x". Then the integral is:


"\int_1^3 f'(x)dx = f(3)-f(1) = 2\cdot 3^2 - 2\cdot 1^2 =16"